OpenStudy (anonymous):

WHERE DO I GO FROM HERE??? I'm trying to get 'w' by itself... 2w+w√2=10

OpenStudy (anonymous):

you can factor the "w" out of the two terms on the left: 2w+w√2=10 w(2 + √2) = 10 Then you can divide both sides by the stuff in parenthesis: w(2 + √2) / (2 + √2) = 10 / (2 + √2) w = 10 / (2 + √2)